# GREEN: Generative Radiology Report Evaluation and Error Notation

### v1

@@ Abstract
Evaluating radiology reports is a challenging problem as factual correctness is extremely important due to the need for accurate medical communication about medical images. Existing automatic evaluation metrics either suffer from failing to consider factual correctness (e.g., BLEU and ROUGE) or are limited in their interpretability (e.g., F1CheXpert and F1RadGraph). In this paper, we introduce GREEN (Generative Radiology Report Evaluation and Error Notation), a radiology report generation metric that leverages the natural language understanding of language models to identify and explain clinically significant errors in candidate reports, both quantitatively and qualitatively. Compared to current metrics, GREEN offers: 1) a score aligned with expert preferences, 2) human interpretable explanations of clinically significant errors, enabling feedback loops with end-users, and 3) a lightweight open-source method that reaches the performance of commercial counterparts. We validate our GREEN metric by comparing it to GPT-4, as well as to error counts of 6 experts and preferences of 2 experts. Our method demonstrates not only higher correlation with expert error counts, but simultaneously higher alignment with expert preferences when compared to previous approaches."
@@ 1 Introduction
Machine learning has enabled great progress in the automatic interpretation of images, where vision language models (VLMs) translate features of images into text Radford et al. (2021); Liu et al. (2024). In the medical domain, patient images are interpreted by radiologists, which is referred to as radiology report generation (RRG). Automated and high-quality RRG has the potential to greatly reduce the repetitive work of radiologists, alleviate burdens arising from shortage of radiologists, generally improve clinical communication Kahn Jr et al. (2009), and increase the accuracy of radiology reports Rajpurkar and Lungren (2023).
@@ 3 GREEN
First, we describe the construction of our generative language model, which is designed to identify and classify errors in radiology reports into six categories (Section 3.1). This section is subdivided into the collection of training data (Section 3.1.1) and details of the training process (Section 3.1.2). Second, we elaborate on the GREEN score in Section 3.2, including its rationale and significance. Third, we explain the GREEN summary and its usefulness.
@@ 3.1 Generative Large Language Model
3.1.1 Chest X-rays Data Collection
To compile our training dataset, we selected 100,000 reference and generated candidate report pairs from six publicly-available de-identified chest X-ray datasets: MIMIC-CXR Johnson et al. (2019), MIMIC-PRO Ramesh et al. (2022), CandidPTX Feng et al. (2021), PadChest Bustos et al. (2020), BIMCV-covid19 Vayá et al. (2020) and OpenI Demner-Fushman et al. (2012). We employed the prompt shown in Appendix A.1 to task GPT-4 to identify and categorize differences in natural language across six unique clinically-defined categories detailed previously Yu et al. (2023b).
The pairing process used five different heuristics, generating 20,000 unique pairs for each heuristic: i) randomly matching candidates and references, ii) modifying the candidate by removing and shuffling sentences from the original report, iii) using a trained RRG model to create the candidate based on the referenced image (to bridge the gap noted in literature between actual reports and those generated by baseline models), iv) pairing candidates with the closest semantically similar report assessed using BERTScore Zhang et al. (2019), and v) creating candidates through RadGraph (named-entity recognition dataset) Jain et al. (2021) permutations of the reference reports, incorporating changes to the presence of findings or by making modifications to size, severity, or location throughout the reports. The number of RadGraph permutations and the BERTScore distribution of the pairs are presented in Figure 3. Additionally, a sample GPT-4 response of a pair with a candidate that includes exactly one RadGraph permutation can be found in Appendix A.2.
@@ 3.1.1 Chest X-rays Data Collection
[Clinically Significant Errors]:
@@
3.1.2 Model Architecture and Training
To enhance performance with medical data, we pre-trained LLaMA-2 and Phi-2 using a comprehensive set of domain-specific datasets to form RadLLaMA-2 and RadPhi-2. These datasets include MIMIC-IV Radiology Reports Johnson et al. (2023), MIMIC-IV Discharge Summaries, MIMIC-CXR Radiology Reports, and a variety of sources from PubMed (Abstracts and Patient Reports). Specialized datasets such as Wiki Medical Terms111www.huggingface.co/datasets/gamino/wiki_medical_terms and Medical Guidelines222www.huggingface.co/datasets/epfl-llm/guidelines Vashishth et al. (2021) were also used.
@@ 3.3 GREEN Summary
We present a method for a detailed analysis of error explanation per error subcategory $i$. Initially, the process involves clustering the embeddings of the explanation sentence, $e_{i}\in E_{i}$, into $k$ clusters using Sentence Transformers Ni et al. (2021). Subsequent steps include quantifying the size of these $k$ clusters, where $k$ is determined by the silhouette distance Shahapure and Nicholas (2020). We then identify the largest cluster and isolate the top-3 members exhibiting the closest distance to the cluster’s mean as determined through cosine distance measures. The top-3 members are then named for each error category in the GREEN summary (Appendix A.5). For example, in Figure 4, subcategory $(a)$ has "(Small right pleural effusion, Small right pleural effusion, Small right pleural effusion)" as the three closest members to the largest cluster, which indicates significant hallucinations of a "small right pleural effusion". This opens up the possibility for targeted detection of data biases or quality issues, as well as specific areas for model improvement.
@@ 3.4.2 Expert Preference Dataset
Since we lacked a ground-truth score for evaluating radiology reports, we turned to a radiologist preference dataset. This helped to address the shortcomings of comparing errors counts in the previous section. In particular, when radiologists compare two reports, they do so with an intuitive weighting of matched findings and significant and insignificant errors. As such, the preference dataset enabled us to determine which metric most effectively replicates expert evaluations and allowed us to assess GREEN as a preference generator  Rafailov et al. (2024); Ethayarajh et al. (2024).
@@ 4.1 Inter-Expert Analysis
The baseline for model performance was established by comparing the correlation between experts to each other from the ReXVal dataset using Kendall’s Tau coefficient Yu et al. (2023b). The correlation between the 6 experts was less than the average correlation across experts, which spans from 0.41 to 0.60 (Figure 5). Additionally, we assessed the discrepancy between experts by computing the mean absolute error of significant error counts, resulting in a mean difference of 0.83 ± 0.13 (Table LABEL:tab:inter_expert_mae). These inter-expert measures serve as upper bound for the GREEN model performance outlined below.
@@ 4.2 Performance on Training Data Distribution
Table 4: Results on the external validation set (200 examples) compared to ReXVal human experts. 1MAE: Mean Absolute Error of the sum of Sig. Errors and Insig. Errors, 2STD: Standard Deviation, 3Average error count. (a)-(f) Accuracy for each significant error category: (a) False report of a finding in the candidate, (b) Missing a finding present in the reference, (c) Misidentification of a finding’s anatomic location/position, (d) Misassessment of the severity of a finding, (e) Mentioning a comparison that isn’t in the reference, and (f) Omitting a comparison detailing a change from a prior study.
Language Model MAE1 ± STD 2 Accuracy $\uparrow$
[…]
7.03 ± 1.16 3 0.47 ± 0.523
Compared to Mistral-v0.1 (7B) 2.60 ± 1.91 0.87 ± 0.94 0.13 0.31 0.62 0.59 0.48 0.67
Average Expert LLaMA-2 (7B) 2.62 ± 1.25 0.47 ± 0.52 0.10 0.23 0.65 0.59 0.68 0.70
Phi-2 (2.7B) 2.10 ± 1.39 0.65 ± 0.70 0.34 0.08 0.65 0.57 0.66 0.53
RadLLaMA-2 (7B) 1.54 ± 1.36 0.51 ± 0.54 0.34 0.38 0.60 0.54 0.65 0.68
RadPhi-2 (7B) 2.08 ± 1.15 0.55 ± 0.61 0.19 0.18 0.62 0.57 0.62 0.61
GPT-4 1.51 ± 1.29 0.52 ± 0.55 0.32 0.40 0.65 0.59 0.68 0.70
@@ 4.3.1 Expert Error Counts
To quantitatively validate GREEN, we measured the mean absolute difference and accuracy relative to the average radiologist, as detailed in Section 4.1. We found that, overall, RadLLaMA-2 exhibits the lowest differences to the mean radiologist’s error counts (1.54 ± 1.36 sig. error difference), which approaches the performance of GPT-4 (1.51 ± 1.29 sig. error difference). Compared to all experts individually, RadLLaMA-2 exhibits an average difference of 1.02 ± 0.27, which is within the boundaries of the average inter-expert difference of 0.83 ± 0.13. Drawing from these quantitative results, along with the findings presented in Section 3.4, we selected RadLLaMA-2 as the GREEN model for all future experiments.
@@
To validate GREEN against existing metrics, we assessed the correlation between the total error count by radiologists and the classical metrics, alongside GPT-4 and GREEN, using an external dataset (Table 5). Both GREEN-GPT-4 and GREEN demonstrated similar and stronger correlations compared to classical metrics. We noted that the GREEN correlation significantly outperforms that of RadGraph, despite RadGraph being trained on human annotations (RadGraph: 0.47 (95% CI, -0.55 0.39) vs. GREEN: 0.63 (95% CI, 0.69 0.56) ).
Compared to the inter-expert correlation, GREEN exhibits a competitive degree of correlation at 0.63 compared to the range from 0.48 to 0.64 on the same examples (Figure 5).
Additionally, we observed that correlating the unweighted summed error counts of GREEN (clinically sig. errors + insig. errors) yields a correlation coefficient of 0.79 (95% CI, 0.74 0.83), which may exceed the performance of GPT-4-based G-Rad Chaves et al. (2024).
@@
Table 5: Correlation analysis between metrics, GREEN score, and GREEN score inferred from GPT-4 (GREEN-GPT-4) to average total error count of 6 radiologists in the ReXVal dataset (200 examples).
@@
BLEU 0.33 (95% CI, 0.42 0.23)
BERTScore 0.44 (95% CI, 0.52 0.35)
F1RadGraph 0.47 (95% CI, 0.55 0.39)
ROUGE-L 0.53 (95% CI, 0.61 0.45)
RadCliQ-v1 0.58 (95% CI, 0.51 0.64)
GREEN (ours) 0.63 (95% CI, 0.69 0.56)
GREEN GPT-4 (ours) 0.64 (95% CI, 0.70 0.57)
Error counts GREEN 0.79 (95% CI, 0.74 0.83)
Error counts GPT-4 0.79 (95% CI, 0.75 0.83)
4.3.2 Expert Preferences
@@
We observed the highest mean accuracy for GREEN and GREEN-GPT-4, which both outperforms the summed error count approach. The preferences of GPT-4 exhibited an accuracy of 0.23 (95% CI, 0.13 0.36) (Table LABEL:tab:preferences).
Upon examining the impact of varying confidence levels (Figure 6), we observed that GREEN’s preference alignment improves in conjunction with increased radiologist confidence, distinguishing it from the approach of using just the sum of error counts or the direct GPT-4 preference with low accuracy.

### v2

@@ Abstract
Evaluating radiology reports is a challenging problem as factual correctness is extremely important due to the need for accurate medical communication about medical images. Existing automatic evaluation metrics either suffer from failing to consider factual correctness (e.g., BLEU and ROUGE) or are limited in their interpretability (e.g., F1CheXpert and F1RadGraph). In this paper, we introduce GREEN (Generative Radiology Report Evaluation and Error Notation), a radiology report generation metric that leverages the natural language understanding of language models to identify and explain clinically significant errors in candidate reports, both quantitatively and qualitatively. Compared to current metrics, GREEN offers: 1) a score aligned with expert preferences, 2) human interpretable explanations of clinically significant errors, enabling feedback loops with end-users, and 3) a lightweight open-source method that reaches the performance of commercial counterparts. We validate our GREEN metric by comparing it to GPT-4, as well as to error counts of 6 experts and preferences of 2 experts. Our method demonstrates not only higher correlation with expert error counts, but simultaneously higher alignment with expert preferences when compared to previous approaches. We publish the code as a pypi package and datsets  [^1] https://stanford-aimi.github.io/green.html.
@@ 1 Introduction
Machine learning has enabled great progress in the automatic interpretation of images, where vision language models (VLMs) translate features of images into text Radford et al. (2021); Liu et al. (2024). In the medical domain, patient images are interpreted by radiologists, which is referred to as radiology report generation (RRG). Automated and high-quality RRG has the potential to greatly reduce the workload of radiologists alleviating burdens arising from shortage of radiologists, generally improve clinical communication Kahn Jr et al. (2009), and increase the accuracy of radiology reports Rajpurkar and Lungren (2023).
@@ 3 GREEN
First, we describe the construction of our generative language model, which is designed to identify and classify errors in radiology reports into six categories (Section 3.1). This section is subdivided into a process overview, collection of the reference and synthetic candidate reports (Section 3.1.1) and details of the training process (Section 3.1.2). Second, we elaborate on the GREEN score in Section 3.2, including its rationale and significance. Third, we explain the text-form GREEN summary and its usefulness.
@@ 3.1 Generative Large Language Model
We designed a prompt for GPT-4 and LLM-fintuning consisting of the a process overview, criteria for judgement (six distinct error types for each clinically significant and insignificant errors), a reference and a synthetically modified candidate report and a response template (See prompt at A.1 and response at A.2, Figure 2). Second, we generated a dataset by prompting GPT-4 with the prompt and received a response, called GREEN analysis. Lastely, we used the GREEN analysis to obtain and validate the GREEN score (numerical) and GREEN summary (free text).
3.1.1 Reference and Synthetic Candidate Reports
To compile pairs of reference and candidate reports, we selected 100,000 reference and generated candidate report pairs from six publicly-available de-identified chest X-ray datasets: MIMIC-CXR Johnson et al. (2019), MIMIC-PRO Ramesh et al. (2022), CandidPTX Feng et al. (2021), PadChest Bustos et al. (2020), BIMCV-covid19 Vayá et al. (2020) and OpenI Demner-Fushman et al. (2012). We employed the prompt shown in Appendix A.1 to task GPT-4 to identify and categorize differences in natural language across six unique clinically-defined categories detailed previously Yu et al. (2023b).
The pairing process used five different heuristics, generating 20,000 unique pairs for each heuristic: i) randomly matching candidates and references, ii) modifying the candidate by removing and shuffling sentences from the original report, iii) using a trained RRG model to create the candidate based on the referenced image, iv) pairing candidates with the closest semantically similar report assessed using BERTScore Zhang et al. (2019), and v) creating candidates through RadGraph (named-entity recognition dataset) Jain et al. (2021) permutations of the reference reports, incorporating changes to the presence of findings or by making modifications to size, severity, or location throughout the reports. The number of RadGraph permutations and the BERTScore distribution of the pairs are presented in Figure 3. Additionally, a sample GPT-4 response of a pair with a candidate that includes exactly one RadGraph permutation can be found in Appendix A.2.
@@ 3.1.1 Reference and Synthetic Candidate Reports
[Clinically Significant Errors]: <accuracy>
<most representative error>
@@
3.1.2 Baseline Model and Training
To enhance performance with medical data, we pre-trained LLaMA-2 and Phi-2 using a comprehensive set of domain-specific datasets to form RadLLaMA-2 and RadPhi-2. These datasets include MIMIC-IV Radiology Reports Johnson et al. (2023), MIMIC-IV Discharge Summaries, MIMIC-CXR Radiology Reports, and a variety of sources from PubMed (Abstracts and Patient Reports). Specialized datasets such as Wiki Medical Terms222www.huggingface.co/datasets/gamino/wiki_medical_terms and Medical Guidelines333www.huggingface.co/datasets/epfl-llm/guidelines Vashishth et al. (2021) were also used.
@@ 3.2 GREEN Score
The GREEN score rewards matched findings, balancing the penalties for clinically significant errors by employing an inverse structure. Consequently, as reports become more concise and accurate, achieving a higher score becomes increasingly challenging.
@@ 3.3 GREEN Summary
To the best of the authors knowledge, we present the first method for a detailed free text analysis of error explanation per error subcategory for Chest-Xray report generation. The GREEN analysis (model response) contains three parts: the explanation, clinically significant and insignificant error counts, matched findings. We use clinically significant error counts which would alter the clinical workflow. This part is further divided into six subcategories. The model response for each subcategory includes the error counts and an explanation sentence about each error. We gather these explanation sentences, $e_{i}\in E_{i}$, for each subcategory $i$ and embed them using Sentence Transformers Ni et al. (2021). The embeddings are then grouped into $k$ cluster, where $k$ is determined by the silhouette distance Shahapure and Nicholas (2020). We choose the largest cluster and the closest to the mean embeddings of the largest cluster. We map these embeddings back to the respective individual sentences, which are then included in the GREEN summary (Figure 4). For example, in Figure 4, subcategory $(a)$ has "(Small right pleural effusion, Small right pleural effusion, Small right pleural effusion)" as the three closest members to the largest cluster, which indicates significant hallucinations of a "small right pleural effusion". This opens up the possibility for targeted detection of data biases or quality issues, as well as specific areas for model improvement.
@@ 3.4.2 Expert Preference Dataset
Since we lack a ground-truth score for evaluating radiology reports, we turned to a radiologist preference dataset. This helped to address the shortcomings of just comparing errors counts in the previous section, may reduce score overfitting to the RexVal dataset. In particular, when radiologists compare two reports, they do so with an intuitive weighting of matched findings and significant and insignificant errors. As such, the preference dataset enabled us to determine which metric most effectively replicates expert evaluations and allowed us to assess GREEN as a preference generator  Rafailov et al. (2024); Ethayarajh et al. (2024).
@@ 4.1 Inter-Expert Analysis
The baseline for model performance was established by comparing the correlation between experts to each other from the ReXVal dataset using Kendall’s Tau coefficient Yu et al. (2023b). The correlation between the 6 experts was less than the average correlation across experts, which spans from 0.41 to 0.60 (Appendix Figure 6). Additionally, we assessed the discrepancy between experts by computing the mean absolute error of significant error counts, resulting in a mean difference of 0.83 ± 0.13 (Table LABEL:tab:inter_expert_mae). These inter-expert measures serve as upper bound for the GREEN model performance outlined below. For testing the difference in location of the mean difference between expert, we calculate the statistics using the Wilcoxon-signed rank test. These values show that the experts themselves already significantly differ, leading us to the conclusion that the Wilcoxon-signed rank test is less suitable for comparison of error counts between GREEN, GPT, and experts (Appendix Tables 10 and 11).
@@ 4.3.1 Expert Error Counts
To quantitatively validate GREEN, we measured the mean absolute difference and accuracy relative to the average radiologist, as detailed in Section 4.1. We found that, overall, RadLLaMA-2 exhibits the lowest differences to the mean radiologist’s error counts (1.54 ± 1.36 sig. error difference), which approaches the performance of GPT-4 (1.51 ± 1.29 sig. error difference). Compared to all experts individually, RadLLaMA-2 exhibits an average difference of 1.02 ± 0.27, which is within the boundaries of the average inter-expert difference of 0.83 ± 0.13. Drawing from these quantitative results, along with the findings presented in Section 3.4, we selected RadLLaMA-2 as the GREEN model for all future experiments (Table
Table 4: Results on the external validation set (200 examples) compared to ReXVal human experts. 1MAE: Mean Absolute Error of the sum of Sig. Errors and Insig. Errors, 2STD: Standard Deviation, 3Average error count. (a)-(f) Accuracy for each significant error category: (a) False report of a finding in the candidate, (b) Missing a finding present in the reference, (c) Misidentification of a finding’s anatomic location/position, (d) Misassessment of the severity of a finding, (e) Mentioning a comparison that isn’t in the reference, and (f) Omitting a comparison detailing a change from a prior study.4PreRad=pretrained on radiology text, Section 3.1.2, 5MM=multimodal, described in Section 5.1.
Language Model Data MAE1 ± STD 2 Accuracy $\uparrow$
[…]
PreRad4 CXR MM5 7.03 ± 1.16 3 0.47 ± 0.523
Mistral-v0.1 (7B) 2.60 ± 1.91 0.87 ± 0.94 0.13 0.31 0.62 0.59 0.48 0.67
LLaMA-2 (7B) 2.62 ± 1.25 0.47 ± 0.52 0.10 0.23 0.65 0.59 0.68 0.70
LLaMA-2 (7B) 1.54 ± 1.36 0.51 ± 0.54 0.34 0.38 0.60 0.54 0.65 0.68
Phi-2 (2.7B) 2.10 ± 1.39 0.65 ± 0.70 0.34 0.08 0.65 0.57 0.66 0.53
Phi-2 (2.7B) 2.08 ± 1.15 0.55 ± 0.61 0.19 0.18 0.62 0.57 0.62 0.61
GREEN GPT-4 1.51 ± 1.29 0.52 ± 0.55 0.32 0.40 0.65 0.59 0.68 0.70
@@
To validate GREEN against existing metrics, we assessed the correlation between the radiologists total error count (sum of significant and insignificant errors) and the classical metrics, alongside GPT-4 and GREEN as proposed in Yu et al. (2023b). Both GREEN-GPT-4 and GREEN demonstrated similar and stronger correlations compared to classical metrics. We noted that the GREEN correlation significantly outperforms that of RadGraph, despite RadGraph being trained on human annotations (RadGraph: 0.47 (95% CI, -0.55 0.39) vs. GREEN: 0.63 (95% CI, 0.69 0.56) )(Table 6).
Compared to the inter-expert correlation, GREEN exhibits a competitive degree of correlation at 0.63 compared to the range from 0.48 to 0.64 on the same examples (Figure 6). While this approach enables cross-metric and cross-study comparisons, it may not be clinically optimal. The total error count fails to differentiate between error significance, compares scores on different scales (0 to 1 versus infinity to 0) and newly proposed scores may overfit to this benchmark design.
Furthermore, our analysis revealed that the correlation coefficient between the unweighted total error count of GREEN’s (summing clinically significant and insignificant errors) and radiologists’ total error counts is 0.79 (95% CI: 0.74-0.83). This correlation is statistically indistinguishable from the experts total error and the performance of both GPT-4-based G-Rad  Chaves et al. (2024) and GPT-4-based GREEN when compared to radiologists’ total error counts.
For more fine-grained analyses we calculate the error count differences for each subcategory (Table 5). The GREEN model only has one shaded cell for category C, indicating that the largest deviations are “Misidentification of a finding’s anatomic location/position” compared to all 6 experts and GPT-4.
Table 5: Fine grained error difference to the mean rater, subcategories (a)-(f): Accuracy for each significant error category: (a) False report of a finding in the candidate, (b) Missing a finding present in the reference, (c) Misidentification of a finding’s anatomic location/position, (d) Misassessment of the severity of a finding, (e) Mentioning a comparison that isn’t in the reference, and (f) Omitting a comparison detailing a change from a prior study.
Rater (a) (b) (c) (d) (e) (f)
Expert 0 $0.28\pm 0.38$ $0.26\pm 0.37$ $0.15\pm 0.26$ $0.12\pm 0.21$ $0.11\pm 0.21$ $0.09\pm 0.17$
Expert 1 $0.30\pm 0.37$ $0.25\pm 0.38$ $0.15\pm 0.26$ $0.13\pm 0.23$ $0.10\pm 0.18$ $0.09\pm 0.17$
Expert 2 $0.30\pm 0.36$ $0.26\pm 0.46$ $0.16\pm 0.29$ $0.24\pm 0.40$ $0.10\pm 0.19$ $0.10\pm 0.20$
Expert 3 $0.28\pm 0.36$ $0.28\pm 0.39$ $0.14\pm 0.25$ $0.15\pm 0.27$ $0.18\pm 0.37$ $0.18\pm 0.36$
Expert 4 $0.34\pm 0.42$ $0.22\pm 0.32$ $0.16\pm 0.29$ $0.16\pm 0.26$ $0.15\pm 0.29$ $0.14\pm 0.27$
Expert 5 $0.67\pm 0.77$ $0.37\pm 0.55$ $0.13\pm 0.23$ $0.19\pm 0.34$ $0.12\pm 0.22$ $0.08\pm 0.15$
GREEN 0.36 ± 0.45 0.38 ± 0.48 0.21 ± 0.34 0.23 ± 0.36 0.14 ± 0.25 0.11 ± 0.21
GREEN GPT-4 0.32 ± 0.37 0.43 ± 0.54 0.15 ± 0.28 0.2 ± 0.3 0.13 ± 0.24 0.14 ± 0.28
4.3.2 Expert Preferences
@@
Table 6: Correlation analysis between metrics to total error count of 6 radiologists in the ReXVal dataset (200 examples).
@@
QAFactEval 0.32 (95% CI, 0.21 0.41)
AlignScore 0.32 (95% CI, 0.23 0.41)
BLEU 0.35 (95% CI, 0.25 0.43)
BERTScore 0.49 (95% CI, 0.40 0.56)
F1RadGraph 0.57 (95% CI, 0.49 0.63)
ROUGE-L 0.56 (95% CI, 0.48 0.63)
RadCliQ-v1 0.61 (95% CI, 0.55 0.66)
GREEN (ours) 0.63 (95% CI, 0.56 0.69)
GREEN GPT-4 (ours) 0.64 (95% CI, 0.57 0.70)
Error count G-Rad (GPT-4) 0.76 (95% CI, 0.70, 0.80)
Error count GREEN 0.79 (95% CI, 0.74 0.83)
Error count GPT-4 0.79 (95% CI, 0.75 0.83)
Inter-Expert 0.81 (95% CI, 0.78 0.83)
@@
We observed the highest accuracy for GREEN and GREEN-GPT-4, which both outperforms the summed error count approach. The preferences of GPT-4 exhibited an accuracy of 0.23 (95% CI, 0.13 0.36) (Table LABEL:tab:preferences).
Upon examining the impact of varying confidence levels (Figure 5), we observed that GREEN’s preference alignment improves in conjunction with increased radiologist confidence, distinguishing it from the approach of using just the total error counts or the direct GPT-4 preference with low accuracy.